OpenStudy (anonymous):

Give the 3rd and 7th terms using the sequence formula tn = (n – 3)2

OpenStudy (anonymous):

I haven't done much for sequences, but I think I can solve this if I understand the notation: Does this read as, t of n = (n-3) squared?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Assuming that the sequence starts at n=1, the third term is \[t_{3} = (3-3)^2=0^2=0\] and the 7th term is \[t_{7} = (7-3)^2=4^2=16\]

OpenStudy (anonymous):

Aren't we also supposed to plug in the value of the previous function? Or am I mixing it up with summation?

jimthompson5910 (jim_thompson5910):

You're thinking of a recursive sequence uberl337h4xx0r, and summation generally refers to a series.